### v1

# Let’s Get to the Point: LLM-Supported Planning, Drafting, and Revising of Research-Paper Blog Posts

Abstract.
Research-paper blog posts help scientists disseminate their work to a larger audience, but translating papers into this format requires substantial additional effort. Blog post creation is not simply transforming a long-form article into a short output, as studied in most prior work on human-AI summarization. In contrast, blog posts are typically full-length articles that require a combination of strategic planning grounded in the source document, well-organized drafting, and thoughtful revisions. Can tools powered by large language models (LLMs) assist scientists in writing research-paper blog posts? To investigate this question, we conducted a formative study (N=6) to understand the main challenges of writing such blog posts with an LLM: high interaction costs for 1) reviewing and utilizing the paper content and 2) recurrent sub-tasks of generating and modifying the long-form output. To address these challenges, we developed Papers-to-Posts, an LLM-powered tool that implements a new Plan-Draft-Revise workflow, which 1) leverages an LLM to generate bullet points from the full paper to help users find and select content to include (Plan) and 2) provides default yet customizable LLM instructions for generating and modifying text (Draft, Revise). Through a within-subjects lab study (N=20) and between-subjects deployment study (N=37 blog posts, 26 participants) in which participants wrote blog posts about their papers, we compared Papers-to-Posts to a strong baseline tool that provides an LLM-generated draft and access to free-form LLM prompting. Results show that Papers-to-Posts helped researchers to 1) write significantly more satisfying blog posts and make significantly more changes to their blog posts in a fixed amount of time without a significant change in cognitive load (lab) and 2) make more changes to their blog posts for a fixed number of writing actions (deployment).
1. Introduction
Sharing scientific ideas outside of research papers is important for increasing awareness and impact (Jucan and Jucan, 2014; Burns et al., 2003). Scientists often seek different ways to communicate their work to broader audiences, such as through tweets or blog posts (Gero et al., 2021; Long et al., 2023; Williams et al., 2022; Holmberg and Thelwall, 2014); indeed, tweeting about research papers has been shown to raise their citation count (Luc et al., 2021). While some tools have been created to help author short-form derivatives like tweets (Long et al., 2023), significant challenges remain for scientists seeking to create longer-form derivatives like blog posts, which require more extensive planning, drafting, and revising writing processes (Flower and Hayes, 1981). In the planning step, the blog post author must review the paper, itself a long document, to find and select several pieces of information to include in the blog post. In the drafting and revising steps, there are many sub-tasks for completing the full blog post, such as creating an engaging hook or reporting results with sufficient detail but in simpler terms than the paper.
Integrating large language models (LLMs) in writing tools offers a promising avenue for improving writing outcomes; however, existing approaches are not designed to help create long-form derivatives like blog posts from long documents like academic papers. Prior work in human-LLM text summarization has focused on supporting short-form summaries, a sentence or paragraph in length (Zhang et al., 2023b; Lai et al., 2022). In the broader area of human-LLM writing, prior work has looked into providing assistance for writing long-form creative texts such as stories (Chung et al., 2022; Goldfarb-Tarrant et al., 2019; Singh et al., 2022) and screenplays (Mirowski et al., 2023), but these tasks do not take a long-form text input. Prior work in human-LLM scientific writing has investigated how LLM tools can assist scientists in writing short-form texts such as tweetorials (Long et al., 2023) and extended metaphors (Kim et al., 2023), but to the best of our knowledge, no work has sought to develop an LLM tool for creating long-form scientific texts, such as research-paper blog posts.
Still, prior work has shown that LLM tools can support planning, drafting, and revising for different forms of writing, such as creative stories (Goldfarb-Tarrant et al., 2019), short summaries (Zhang et al., 2023b), and short scientific texts (Long et al., 2023). Thus, in this paper we ask: how best can an LLM tool provide assistance in dealing with the challenges of writing long-form summaries of scientific long documents? We designed a technology probe of a simple writing interface with preliminary LLM-based features to better understand the costs and benefits of interacting with LLMs for the purpose of planning, drafting, and revising a research-paper blog post. The technology probe provides the paper text, an LLM-generated bullet point outline for the paper, a drafting space to write the blog post in sections, and access to an LLM prompting interface for writing tasks. Through a formative study (N=6), we gathered feedback on how users wish to leverage LLMs in this context and common interaction challenges. For one, the participants valued having a paper outline for finding and selecting relevant information from the paper to include in the blog post, but they struggled with understanding and utilizing it (e.g., understanding how the bullet points were connected to the paper text). Also, participants desired to provide the LLM recurring instructions for generating and modifying text that were not well-supported by the tool (e.g., rewrite a paragraph to better capture readers’ attention, continue writing from a particular sentence).
Based on the identified challenges, we iterated on the technology probe to develop the prototype tool Papers-to-Posts for LLM-assisted paper blog post writing. The tool follows our Plan-Draft-Revise workflow (Figure 1). In the Plan phase, the user may control text generation by selecting a subset of LLM-generated bullet points corresponding to key points in the source; Papers-to-Posts is the first LLM-powered long-document summarization tool to support direct control of content selection through a source outline, maintaining the source’s narrative structure. In the Draft phase, the user may supply additional instructions and preview LLM generations that use the selected bullet points, instructions, and blog post text thus far. In the Revise phase, they may use an LLM to further refine text with macros that encapsulate frequent editing transformations (e.g., to have less jargon, to be more dramatic).
We evaluated Papers-to-Posts through two studies in which participants wrote blog posts for their own papers, and we compared our tool to a baseline tool that provided an LLM-generated draft and integrated prompting. In a within-subjects lab study (N=20) in which researchers wrote blog posts for two of their own papers, one per tool, we found that, by helping authors to plan, draft, and revise their blog post, Papers-to-Posts led to significantly higher writer satisfaction with the blog post output alongside significantly greater change to the provided draft, without increasing task completion time or cognitive load. To study Papers-to-Posts in a more realistic setting, we also conducted an unmonitored, between-subjects deployment study (N=37), in which participants using Papers-to-Posts again made greater change to the provided drafts for a given amount of time spent or number of writing actions. Furthermore, through quantitative survey ratings and behavioral analysis as well as qualitative short answers and semi-structured interviews across the two studies, we observe that an LLM-generated paper outline and affordances for streamlining common long-form text generation and modification sub-tasks contributed to this increased satisfaction. In discussing our results, we also reflect on design implications for other human-LLM scientific long-document summarization tools with long-form output.
In summary, we make the following main contributions:
A formative study (N=6) producing two primary design goals for an LLM-powered tool for writing research-paper blog posts: (1) help users review and select content from a long source document during the planning stage, and (2) transparently streamline the process of instructing the LLM to generate and modify text during the drafting and revising stages.
Papers-to-Posts, a prototype tool for writing research-paper blog posts, which addresses the design goals identified in the formative study and implements the Plan-Draft-Revise workflow (Figure 1).
A within-subjects lab study (N=20 participants) providing evidence that Papers-to-Posts improves upon a powerful baseline (LLM draft and free-form prompting). Papers-to-Posts led to significantly higher writer satisfaction with the final blog post (with comparable task completion time and cognitive load) alongside significantly greater changes to the initial draft.
A field deployment study (N=37 blog posts, 26 participants) comparing Papers-to-Posts to the same strong baseline show that the same benefits hold in realistic scenarios where participants were writing blog posts on their own and sharing them publicly at the end of the study.
Design implications for other human-LLM scientific long-document summarization tools with long-form output: avoid solely post-editing interactions, present long-document outlines to support content selection, and include transparent affordances to streamline sub-tasks in generating and modifying long-form output text.
2. Related Work
2.1. Human-LLM Summarization
Although there is much work on automatic text summarization (Koh et al., 2022; El-Kassas et al., 2021), issues remain with the output summaries, such as inaccuracies and style misalignment (Lu et al., 2023; Koh et al., 2022). To address these issues, other works have looked into having a human and LLM work together to generate a summary (Lai et al., 2022; Moramarco et al., 2021). According to Cheng et al., there are five common methods of human-AI summarization: guiding AI output, selecting or rating AI output, post-editing of AI-generated summaries, interactive editing, and writing with AI assistance (Cheng et al., 2022). Our tool Papers-to-Posts provides affordances related to guiding and post-editing AI output.
Prior work in human-AI summarization has mainly focused on direct model guidance (e.g., utilizing concepts/queries (Zhang et al., 2023b; Bayatmakou et al., 2022; Avinesh and Meyer, 2017; Avinesh et al., 2018; Ding et al., 2023; Wang et al., 2023b), typing (Xie et al., 2023), ratings (Gao et al., 2020; Bohn and Ling, 2021), gaze (Bohn and Ling, 2021)) or simply post-editing (e.g., (Lai et al., 2022; Moramarco et al., 2021)). A few works have supported summarization based on user-selected points from the source (Bayatmakou et al., 2022; Yan et al., 2011; Zhang et al., 2011; Wang et al., 2023a, b), but the points are ranked or clustered by the system without the same narrative structure as the source document. Another work proposed but did not implement or evaluate model guidance through in-context content selection (Gehrmann et al., 2019). Our approach allows model guidance directly grounded in the input source while maintaining the source narrative structure.
Providing more direct control of input content is useful when the writer has significant knowledge of the source and can use reminders of its contents to select what to include in the summary. We focus on summarizing one’s own content, unlike most prior work, with the exception of work on summarizing medical notes that used the post-editing technique (Moramarco et al., 2021; Cai et al., 2022). Thus, we take advantage of the writer’s knowledge of the source document by having them select content in context of the source’s layout.
There are several works on human-LLM summarization of long documents (Zhang et al., 2011; Gehrmann et al., 2019) and some for scientific long documents in particular (Zarinbal et al., 2019; Bayatmakou et al., 2022; Zhang et al., 2023b). Long documents provide a unique challenge for human-LLM summarization, as there is a great deal of information for the human to digest. As writing research-paper blog posts entails summarizing portions of a research paper, this work seeks to support human-LLM long-document summarization.
2.2. Human-LLM Writing
While LLMs have impressive capabilities with respect to quickly generating large amounts of text, interacting with swaths of text can be difficult to manage without additional affordances. Popular chatbot LLMs such as chatGPT (OpenAI, [n. d.]) and BARD (Google, [n. d.]) offer some options for the user to guide their output beyond general prompting. Both provide example inputs for the user to select. chatGPT includes the option to provide custom instructions for the questions ”What would you like chatGPT to know about you to provide better responses?” and ”How would you like chatGPT to respond?”. BARD allows the user to modify its generated text to be shorter, longer, simpler, more casual, more professional, or modified based on a custom prompt. Tools from prior research also provide users with means to guide LLM writing beyond general prompting, such as keyword specifications (Zhang et al., 2023b), style specifications (Ghazvininejad et al., 2017; Coenen et al., 2021), and input modification (Zhang et al., 2011).
With respect to planning text, a unique aspect of our work is providing an LLM-generated outline for source content selection. There are a couple works in human-LLM writing that have utilized outlines, but they focused on outlining output content rather than input content. Zhang et al. introduce the tool VISAR, a human-AI argumentative writing assistant that allows users to iterate on their draft with adjustments to a visual outline of the draft (Zhang et al., 2023a). Dang et al. introduce a text editor that generates summaries of each paragraph written in the editor to comprise a reverse outline of the writing (Dang et al., 2022). The summaries then become affordances to help writers revise their writing (e.g., restructuring the writing). Our work also generates a reverse outline but as an affordance to guide summary generation based on the input content. For summarizing a long source document, an outline is particularly useful, as reviewing the entire source may require substantial time and effort.
With respect to revising text, like BARD (Google, [n. d.]) and other LLM-powered writing tools (NotionAI, [n. d.]), our final tool allows writers to utilize preset modifications, but some of these are set to be more specific to blog post writing. Closely related to our work, Laban et al. present an editing interface to help users revise their writing with LLM edit suggestions that can be integrated into their writing (Laban et al., 2023). This editing interface focuses solely on the revising step and on smaller edits than our tool.
2.3. Human-LLM Scientific Writing
Several works have investigated how to support human-LLM writing. Most of these works focus on tools for creative writing (Singh et al., 2022; Chung et al., 2022; Mirowski et al., 2023), but a few cater to scientific writing. A few works have investigated human-LLM scientific writing for broader science communication tasks that do not involve a source document. Gero et al. present a system for generating scientific ”sparks”, or inspiring sentences (Gero et al., 2022). Long et al. examine how LLM scaffolding can help people generate relatable hooks for complex scientific topics (Long et al., 2023), and Kim et al. investigate how science writers can generate extended metaphors for scientific ideas with the help of an LLM (Kim et al., 2023). The couple works that have explored human-AI scientific summarization do not allow for in-context selection of source content but rely on selection of concepts (Zhang et al., 2023b) or out-of-context sentences extracted based on queries (Bayatmakou et al., 2022).
These previously studied forms of human-LLM scientific writing are the length of a sentence or paragraph. Based on prior work (Mehlenbacher, 2019), popular resources (Kudos, [n. d.]; UWaterloo, [n. d.]; Patrick Dunleavy, [n. d.]; CMU, [n. d.]; Medium, [n. d.]), and feedback from three communications experts, we determined a set of guidelines for writing research-paper blog posts, which indicate that the shortest blog posts are recommended to be around 500 words in length (Medium, [n. d.]; Kudos, [n. d.]; Patrick Dunleavy, [n. d.]), making blog posts a long-form output. Because research-paper blog posts are longer than a sentence or a paragraph, they require coordinating more writing sub-tasks, such as rewriting a sentence about a method to be easier for a broader audience to understand or summarizing a paper’s results section.
2.4. Scientific Blog Posts
Scientists can engage with broad community through various channels, including press releases, magazines, journals, and tweets (August et al., 2020; Gero et al., 2022). This work focuses on scientific blog posts, which offers increased recognition and transparency of one’s work, writing practice, editorial freedom, and public engagement (Jarreau, 2015; Blanchard, 2011; Zou and Hyland, 2019). There are several forms of scientific blog posts, including academic commentary, free commentary, political commentary, and mediation of research to laypersons (Mahrt and Puschmann, 2014). Here, we focus on blog posts with the primary goal of communicating information from a research paper to scientists in the broad area of the paper (e.g., computer science, psychology).
3. Formative Study and System Iteration
We conducted a formative study to identify challenges that people encounter in working with an LLM-powered technology probe for writing blog posts about their research papers, and corresponding design goals for our final LLM tool.
3.1. Participants
Via academic social networks, we recruited 6 participants for the formative study and compensated them with $25 for the 45-minute study. Participants were PhD students, a population frequently tasked with writing blog posts for papers they have authored. Five worked in computer science and one in information science. To circumvent initial time required for getting used to LLMs themselves, we recruited people who already had some experience of working with LLMs in general (6-10 sittings: 1; ¿10 sittings: 5). We also required participants to have previously written research papers (3-10 papers: 3; ¿10 papers: 3), since the purpose of such a tool is to write blog posts for one’s own research papers. Participants had a mixture of experience with writing research-paper blog posts (0 posts: 4; 1-2 posts: 2). Also, for sufficient participant motivation, we made sure that all participants were more interested than not (at least 5 on a 7-point Likert scale where 7 indicates the highest interest) in writing blog posts for their research papers.
3.2. Study Procedure
The study consisted of 3 iterations, each of which involved sessions with 2 participants. Based on participants’ interactions and feedback, we updated the probe after each iteration for better insight into remaining challenges. Prior to each study session, the participant completed a participation agreement. The session coordinator recorded and transcribed each session in Google Hangouts. The coordinator asked the participant to assume their audience was laypeople, as a common advantage of writing scientific blog posts is to communicate with a broader audience than the paper does (Jarreau, 2015; Blanchard, 2011). The participant then used a link to access the technology probe initialized for one of their research papers that they provided. In order to see how participants would intuitively use the probe, the coordinator purposefully did not initially explain how to use it. For 30 minutes, the participant used the probe to write as much as possible for a blog post about their research paper. We asked participants to think aloud as they worked (Van Someren et al., 1994). The coordinator occasionally asked follow-up questions about what the participant said aloud. Participants could ask for help when they did not understand an aspect of the tool. Once 30 minutes had passed, the coordinator engaged the participant in a semi-structured interview. They asked the participant about the difficulties and benefits of using the probe to 1) navigate the paper and outline, 2) generate and edit text for a blog post section, and 3) edit the blog post layout. Finally, they asked for the participant’s thoughts on receiving different types of feedback on their blog post, but this topic was left for future work.
3.3. Initial Technology Probe
In the first iteration of the technology probe interface (Appendix 28), the left side consisted of blog post sections pre-populated with an LLM-generated draft for drafting and revising the blog post. The right side contained the paper text with selectable paragraphs and an LLM-generated paper outline with bullet points for planning. Under the currently selected blog post section, there was an LLM-supported text generator that contained buttons to generate section text from the selected paragraphs and bullet points as well as to expand and condense the text in the blog post section editor.
3.4. Design Goals and Tool Iteration
After each iteration of the formative study, the first author conducted inductive thematic analysis (Braun and Clarke, 2006) of the participants’ think-aloud and interview comments on their experience with the probe. Participants described difficulties in working with the probe that, aside from general usability issues, may be classified under two overarching challenges.
3.4.1. Challenge 1: Planning Paper Content to Use
An academic paper is a long document with dense information, so we expected that identifying information to include in the blog post would be a challenge. For this reason, our technology probe parsed the paper text and generated a paper outline with an LLM in order to support navigating and utilizing the paper. Still, participants faced difficulties with finding and selecting information from the paper. Our associated design goal was to help writers plan their writing by facilitating review of a long document’s contents and selection of essential information to include in the writing (DG1).
Although participants faced issues with the paper outline, four participants expressed that they liked having the paper outline, and four appreciated the ability to select paper paragraphs or bullet points from which to generate text. Because participants overall liked and used the paper outline for identifying and using essential information from the paper, we iterated on this affordance to address DG1. We identified the following requested functionality associated with this design goal:
Transparent Connection Between Paragraphs and Bullet Points. In the first iteration, P2 noted confusion around how the bullet points in the outline corresponded to the original paper text, despite the existence of a button to the left of each set of bullet points that revealed the associated paragraph when clicked. We therefore added a bright line to the left of each set of bullet points to better visualize its grouping and its connection to the corresponding paragraph (see Figure 2). In the second iteration, P3 expressed uncertainty about how the selection of paragraphs and bullet points were related. For example, they were surprised that a paragraph was selected but only one of three corresponding bullet points was selected. During the formative study, the selection logic was as follows. If a bullet point is selected, the associated paragraph needs to be selected, but if a paragraph is selected, no bullet point needs to be selected. This was done in an attempt to make it clearer that the associated paragraph was given to the model as context in addition to any bullet points selected. However, this logic seemed to confuse the participants.
Filter for Selected Content. In the third iteration, P6 said that they were not sure if they had selected everything that they wanted to select. Throughout all the iterations, the probe only communicated the number of selected paragraphs and bullet points. P6 desired a way to view only the selected paragraphs and bullet points.
3.4.2. Challenge 2: Drafting and Revising with Recurring LLM Instructions
The second challenge was attempting to provide the LLM instructions for generating and modifying text that were not well-supported by the probe. Some of these instructions were frequently desired, but the participant had no reasonable way to input them or, once directly prompting the LLM was available, had to determine a prompt to fit their needs. Participants also faced confusion regarding the prompts being used in the back-end. Our corresponding design goal was therefore to help writers draft and revise their writing by transparently streamlining commonly desired instructions to the LLM for generating and modifying text (DG2). We identified the following requested functionalities under this design goal:
Transparent Prompt Input. In the second iteration, P3 liked a default blog post section and was confused that it appeared significantly different from the new version they generated without adjusting the pre-selected bullet points. For one, there was a difference in the requested length in the prompt for the initial-draft section versus the live-generation section. Also, the participant was not aware of the section-specific instructions that were provided to the model for the initial-draft section. To provide access to default settings, which may often be desirable, we updated the probe to show these section-specific instructions as preset (but adjustable) custom instructions. While the model was always prompted to add 1 to 2 sentences in the third iteration, P5 was confused regarding if and how much the text would expand after clicking the ”expand” button multiple times. Meanwhile, P6 commented that they were unsure if the model had context for the blog post sections above the one for which they were generating text.
Custom Instructions. In the first iteration, our design probe did not support direct custom instructions. Users could only generate or modify the text using the pre-defined operations. P1 tried to work around this limitation by giving specific instructions to the LLM in the blog post section editor, custom bullet points, and custom instructions. The participant seemed unaware that the model was not taking these instructions into account. To provide users with more control over the generated text, we added a field for custom instructions to the generator box.
New Section Generation. In the second iteration, P4 noted wanting to generate entire new sections with the model, perhaps based on the section header.
Flexible Input for Text Modification. In the first iteration, P2 expected and wanted the ”condense” button to iterate on the generated text rather than the editor text, so we updated the ”expand” and ”condense” buttons to do that. However, in later iterations, P4 expressed interest in the model expanding on something they wrote without reference to the paper, and P6 expressed interest in rewriting portions of the original paper in different manners.
Preset Modification Actions. We observed participants attempting to rewrite text in manners that seemed applicable to many research-paper blog posts. For instance, P1 wanted a button to make text more approachable for their audience of laypeople, while P3 rewrote text that seemed ”too lofty” in its claims. For the third iteration, we therefore added buttons for such modifying actions, which included ”simpler terms”, ”less dramatic”, and ”more dramatic.” When the participant clicked the button, the modification would be applied to the last generation.
4. The Papers-to-Posts System
4.1. Paper Outline and Initial Draft Generation
The paper text was extracted from the paper PDF using a public tool for converting scientific PDFs to HTML.111https://papertohtml.org We manually adjusted the parsed HTML so that the section headers were correct, no section or subsection was completely missing, and only the main paper text was included (i.e., no appendix, references, or footnotes). Within the paper text box, participants could click a button to see a paper outline generated for the paper text. To obtain this paper outline, we used an LLM to generate 1 to 3 bullet points to summarize each paragraph in the paper, depending on the length of the paragraph. To generate the initial blog post draft for the tool, all of the paper’s bullet points were given to the model along with the request to select 10 bullet points relevant to the initial four blog post sections: introduction, methods, results, and conclusion. To generate each draft section, the model was provided the selected bullet points, their associated paragraphs, the portion of the blog post that had already been generated, and section-specific guidelines based on prior work (Mehlenbacher, 2019), popular resources (Kudos, [n. d.]; UWaterloo, [n. d.]; Patrick Dunleavy, [n. d.]; CMU, [n. d.]; Medium, [n. d.]), and feedback from three communications experts. We found that this process of generating one draft section at a time with the context of prior sections helped make the draft more coherent while avoiding overwhelming the model with instructions for many sections. For both the initial draft and user generations, the model was provided the associated paragraphs for any bullet points in order to ground the model in the actual paper. We used GPT-4 for all LLM tasks except for extracting bullet points from a paragraph, which we produced using GPT-3.5-Turbo after observing that it produced shorter bullet points that were quicker to read. It was also many times cheaper for this token-intensive generation task.
4.2. User Interface
Developed over the course of the formative study, the Papers-to-Posts system consists of two main panes split vertically. On the left is the pane for writing the blog post (Figure 2a), and on the right is the area for generating and modifying text. The blog post is divided into sections, which can be named, reorganized, and deleted. Each section is connected to a unique LLM state, so the history of all inputs and outputs for text generation and modification are stored separately for each section. The initial blog post draft is generated as described in Section 3.3. The blog post section currently being edited is highlighted in purple to indicate that it is the section for which text is generated. Using the controls on the top of the left pane, users can view the blog post’s word count, a preview of the full blog post, and tips for writing blog posts based on prior work (Mehlenbacher, 2019) and popular resources (Kudos, [n. d.]; UWaterloo, [n. d.]; Patrick Dunleavy, [n. d.]; CMU, [n. d.]; Medium, [n. d.]), as well as feedback from 3 communications experts. There is a button to submit the blog post at the bottom. The writer can also add a new section to the blog post using the plus-sign button next to each blog post section. When the writer clicks this button, they can either create a blank section or generate a section based on the section header they provide. To generate a new section, we prompt GPT-4 to select 10 bullet points for the section based on the provided header and to generate text for the section focused on those bullet points (Appendix A).
The pane for generating and modifying text on the right has three areas. The first is the planning-support area (Figure 2b), which contains the Paper Outline to address DG1. The Paper Outline supports writers in planning their blog post, through review of the paper’s contents and selection of essential information to include in the post. The paper view at the top of the pane offers the writer two views of the paper in two tabs: Paper Text and Paper Outline. Under the Paper Text tab, the writer may view the original paper text, divided into paragraphs. The paper is parsed into collapsible sections. The figures and tables are presented in a separate final section of the paper. Their captions can be copied for pasting in the editor so that the writer knows where they should go in the final blog post. Under the Paper Outline tab, the writer may view the paper outline, divided into sets of bullet points associated with the original text’s paragraphs. The writer may view a bullet point set’s corresponding paragraph by clicking the button to the left, and the reverse is true under the Paper Text tab. The outline is generated, and the pre-selected bullet points for each blog post section are displayed. The writer may select or deselect any bullet points or paragraphs that they want the LLM to use in generating text for the blog post section on which they are working. Selecting a paragraph does not affect the selection of associated bullet points, and vice versa. On the back-end, the model still receives the context of the associated paragraph for each bullet point selected. There are buttons below the paper view to clear all selected bullet points and paragraphs, filter for only selected bullet points and paragraphs, and close or open all the paper sections. When the filter button is activated, if the writer is in the Paper Text tab, if a paragraph was deselected but a corresponding bullet point is still selected, the tool notes that ”A corresponding bulletpoint(s) is still selected,” and vice versa.
Below the planning-support area is the drafting-support area (Figure 2c). There are several inputs to address DG2 by transparently streamlining commonly desired instructions to the LLM for generating text. First, there is a menu to select the desired length for the generated text. The default length is labeled as ”auto,” and the other options are 1 sentence, 1 paragraph, or a few paragraphs. The ”auto” length is the same as for the initial LLM draft sections (around 125 to 250 words) There is next a toggle turned on by default for the LLM to use the context of blog post sections above the one currently being edited. The writer can optionally utilize text fields to provide custom bullet points, custom instructions, and starting text to the LLM. For the initial introduction, methods, results, and conclusion sections of the blog post, the custom instructions field is pre-filled with the section-specific instructions used to generate those sections. The text generator’s ”generate” button utilizes the inputs to generate text for the current blog post section using GPT-4.
The tool also provides a revising-support area (Figure 3d, located below the drafting-support area in the tool), which addresses DG2 by transparently streamlining commonly desired instructions to the LLM for modifying text. First, there is a text field for inputting the text to be modified. This can be anything, including text from the blog post draft, the paper, or a generation. There is then a menu to select the desired length for the modified text. The options are the same as for generated text, with ”auto” being unique to each modifying action (see Appendix A). Next, there is a toggle, on by default, for the LLM to use the context of blog post sections above the one currently being edited. There is also a toggle turned on by default for the LLM to base modified writing on selected paragraphs and bullet points. Underneath the toggles, there is a text field for providing any custom instructions for the modification action. At the bottom, there are 5 preset modification buttons: ”expand”, ”condense”, ”simpler terms”, ”less dramatic”, and ”more dramatic”. There is also a ”custom modification” button. Aside from the ”expand” and ”condense” buttons, all of the affordances for generating and modifying text were additions made to support recurring types of LLM instructions observed in the formative study. In the lab study, participants utilized each affordance to some degree.
5. Lab Study
We conducted a lab study to determine if and how the LLM tool resulting from the formative study may support researchers in writing blog posts about their research papers.
5.1. Hypotheses
We hypothesized that, compared to a baseline consisting of an LLM draft and the ability to prompt the LLM, Papers-to-Posts scaffolding and affordances would provide participants a better experience. Our individual hypotheses were as follows:
H1: Papers-to-Posts leads to higher satisfaction with the output blog post than the baseline.
H2: Papers-to-Posts leads to higher satisfaction with the tool than the baseline.
H3: Papers-to-Posts leads to lower cognitive load than the baseline.
H4: Papers-to-Posts leads to lower task completion time than the baseline in generating a blog post that the writer would be comfortable posting publicly.
5.2. Participants
We recruited 20 participants (M: 14, W: 6) via academic social networks and institutional mailing lists and compensated them with $83.33 USD for 2.5 hours of their time. Participants were predominantly early career researchers (PhD student: 17, master’s student: 1, postdoc: 1, industry researcher: 1), who are often lead authors tasked with writing blog posts. All participants conducted research in computer science. Eight participants noted being in a subarea related to human-computer interaction. Participants had a broad range of experience in terms of prior LLM use (¿10 sittings: 14, 6-10 sittings: 3, 2-5 sittings: 2, ¡2 sittings: 1) and number of research-paper blog posts authored (0 posts: 14, 1-2 posts: 5, 3-10 posts: 1). In order to simulate a naturalistic setting, participants were required to have authored at least 2 research papers for use in the study (2 papers: 1, 3-10 papers: 14, ¿10 papers: 5) and to be more interested than not (at least 5 on a 7-point Likert scale where 7 indicates the highest interest) in writing blog posts for their research papers.
5.3. Study Conditions
The treatment condition involved interacting with Papers-to-Posts (Figure 2), which is described in Section 4. The tool in the baseline condition was designed to be a simplified version of Papers-to-Posts without affordances developed to address the design goals (Figure 4). The baseline provided input paper text, sourced from the PDF and divided into paragraphs and collapsible sections as in Papers-to-Posts. The baseline also provided the ability to prompt GPT-4 with any instructions. A history of generations and their associated instructions was retained for the user’s perusal. Lastly, the baseline had an editor for writing the blog post, which was pre-populated with an LLM draft generated by GPT-4-32k using the entirety of the input paper. GPT-4-32k was used rather than GPT-4 because the context window for GPT-4 was insufficient for longer research papers. The prompt was designed to mirror the prompt for Papers-to-Posts’s initial draft sections, so the resulting draft also had introduction, methods, results, and conclusion sections. However, the prompt did not contain section-specific instructions. The full prompt can be found in Appendix A.
In each condition, the tool was set up for one of the participant’s research papers submitted in the recruitment survey. The paper was different for each condition. Each paper was required to have been published online between 2018 and the study period in 2023, with the participant as an author. Only one paper was published in 2018, and the rest were published during or after 2020. There was no significant difference in publication dates between conditions (Paired-Samples t-Test, t(19)=0.90, p=n.s.). The papers were parsed as described in Section 3.3. The average paper length across conditions differed by less than 10% (baseline: 8452 words, treatment: 7742 words).
5.4. Study Procedure
The study was within-subjects, so each participant completed two study sessions, one in the treatment condition and one in the baseline condition, in randomized order and following a counter-balanced design. The participants’ selected papers were assigned randomly to conditions. Each study session lasted up to about 75 minutes. To provide sufficient flexibility for recruited participants, they can select two study sessions within a week of each other, but the majority of sessions were a day apart (1 day: 13, 2 days: 3, 3 days: 3, 4 days: 1). The condition order was counter-balanced within the group of participants with sessions 1 day apart and within the group of participants with sessions over 1 day apart.
Participants completed a participation agreement before the study. We recorded and transcribed each session in Google Hangouts. To start, the session coordinator introduced participants to the task of writing a blog post with an AI tool. The coordinator then presented a tutorial on the tool. After, participants had 2 minutes to interact with the tool in the context of a sample paper that was not their own, and they could ask questions about the tool. Participants next opened a tool link initialized with one of their research papers. The coordinator asked them to assume that their audience was people in their broad area, which was computer science for all participants. We made this assumption because it was the most common audience for whom potential recruits wanted to write blog posts (52 of 60), according to the recruitment survey. The coordinator instructed participants to work as if they were not there, but to let them know if any questions about the tool arose. They also instructed participants not to use any tools or information outside of the tool provided. We made an exception when participants wanted to view the PDF of their paper due to issues working with its parsing in the tool. This only occurred twice. Lastly, the coordinator asked participants to let them know if they reached a point at which they would feel comfortable publicly publishing the blog post in its current state.
Each participant had up to one hour to work on the blog post. At the 20-minute and 40-minute marks, the coordinator reminded them to let them know if they reached the point of comfort with publication. If and when the participant reached that point, the coordinator asked them to save their draft and continue revising the blog post. After the first four study sessions, we noticed that participants were sometimes done early and did not desire to revise further. Thus, we added a rule so that the participant would let the coordinator know if they reached a point at which they were proud of the blog post and saw no reason to revise further. If and when the participant reached that point, or if an hour had passed, they were asked to submit and moved on to the follow-up survey regarding their experience with the tool. The survey consisted of Likert-type questions regarding perceived efficiency, cognitive load, ease of using the tool in relation to the design goals, satisfaction with the tool, satisfaction with the final blog post, perceived control, and perceived ownership. There were also short answer questions regarding the difficulties and benefits of using the tool in relation to the design goals. If it was the participant’s second session, the survey ended with a few questions directly comparing the two conditions. These were not for analysis but only to help us get a quick sense of the results. However, they were not allowed to change their first-session responses. At the end of the session, the coordinator told participants to close out of the tool and, if it was their first session, not to use the tool before the next session. Each time the participant prompted the LLM, saved, or submitted, a log of the user action along with the full state of the blog post draft was saved for analysis.
5.5. Results
We evaluated our 4 hypotheses (Section 5.1) for statistical significance and found that blog post satisfaction is significantly different between the two conditions, while tool satisfaction, cognitive load, and difference in task completion time are not. We also investigated how well the design goals were fulfilled, and the extent of participants’ editing. We saw evidence that our tool addressed the design goals and increased participants’ satisfaction with control over the tool. In addition, we observed that participants did significantly more editing of the initial LLM draft using Papers-to-Posts (we note that this significance test was not part of our initial hypotheses and is therefore exploratory in nature).
5.5.1. Blog Post Satisfaction
We utilized the Likert responses (see Figure 5) as a measure of participants’ satisfaction with their final blog post under both conditions. Given the Likert nature of the data, we employed a non-parametric test. Furthermore, we used a sign test rather than a Wilcoxon signed rank test, which is normally used to analyze non-parametric within-subjects data for two conditions of one factor, due to the violation of that test’s assumption of a symmetric distribution of the paired differences about the median. We observed that participants were significantly more satisfied with the final blog post generated with Papers-to-Posts (M=6.00,Q1=5.00,Q3=7.00) as compared to the baseline tool (M=5.00,Q1=5.00,Q3=6.00) (Dependent-Samples Sign Test, S=11, p¡.05). The majority of participants (11) were more satisfied in the treatment condition, 7 were equally satisfied in both conditions, and only 2 were more satisfied in the baseline condition.
Inspecting participants’ Likert responses regarding aspects of their blog post satisfaction, we see that participants particularly demonstrated more satisfaction with the content covered in the Papers-to-Posts blog post than the baseline blog post (Figure 5b). The majority of participants (14) were more satisfied with the content under the treatment condition (equally: 2, less: 4). Overall, participants were more satisfied with the blog post structure using Papers-to-Posts than the baseline (Figure 5d). The plurality of participants (9) were more satisfied with the structure under the treatment (equally: 6, less: 5). Participants did not show increased satisfaction with the blog post’s tone under the treatment compared to the baseline (Figure 5c).
Looking at participants’ perceived task performance (Figure 6c), we observe that participants had higher perceived performance with Papers-to-Posts than with the baseline. The plurality of participants (9) thought they were more successful in writing a research-paper blog post with Papers-to-Posts compared to the baseline (equally: 7, less: 4). This further corroborates participants’ blog post satisfaction under the treatment condition.
5.5.2. Tool Satisfaction
For measuring tool satisfaction, we used the 5-question Likert scale on tool acceptance from Kocielnik et al. (Kocielnik et al., 2019) (Figure 5). The 5 responses were averaged to obtain one score. We used the non-parametric sign test to evaluate the data, as it was Likert data and violated the Wilcoxon signed rank test’s assumption of a symmetric distribution of the paired differences about the median. While participants had a higher median satisfaction with Papers-to-Posts (M=5.9,Q1=5.2,Q3=6.5) than the baseline tool (M=5.6,Q1=4.8,Q3=6.0), this difference was not significant (Dependent-Samples Sign Test, S=13, p=n.s.).
In Figure 5e-h, we see that, under the treatment condition, participants were more likely to find the tool helpful, use it if it were available, recommend it, and perceive it as useful for productivity. For each of these questions, at least 4 more participants increased rather than decreased their Likert rating from the baseline to the treatment condition. However, for the question about the tool inducing annoyance, only 2 more participants indicated decreased rather than increased annoyance with Papers-to-Posts.
5.5.3. Cognitive Load
To measure cognitive load, we used the NASA TLX Index questionnaire (Hart and Staveland, 1988) (Figure 6a-d) and summed the responses for the mental demand, temporal demand, effort, and frustration questions (see 5.5.1 for discussion of the responses for performance). We utilized the non-parametric sign test to evaluate the data, as it was Likert data and violated the Wilcoxon signed rank test’s assumption of a symmetric distribution of the paired differences about the median. We saw no significant difference between the baseline and treatment conditions, although participants reported slightly lower cognitive load with Papers-to-Posts (M=13.0,Q1=11.5,Q3=16.3) than the baseline tool (M=14.5,Q1=12.0, Q3=18.3) (Dependent-Samples Sign Test, S=8, p=n.s.). Figure 6a-d presents little difference in mental demand, temporal demand, and frustration between the treatment and baseline conditions. However, with respect to effort, we note a divergence. The majority of participants (11) reported using less effort with Papers-to-Posts than with the baseline tool, whereas only 4 participants reported using more effort.
5.5.4. Task Completion Time
We recorded two metrics for task completion time. One was the amount of time participants needed during the session before they would be comfortable publicly publishing the blog post. The other was the perceived amount of time participants would need after the session before they would be comfortable publicly publishing the blog post. As participants reached that point of comfort during most sessions, we focused on the former metric. Because this data was numeric within-subjects data for two conditions of one factor, we used a paired-samples t-test. Participants worked slightly faster with the baseline tool (M=44.3 minutes,SD=12.5) in comparison to Papers-to-Posts (M=46.0 minutes,SD=13.9), but this difference was not significant (Paired-Samples t-Test, t(19)=0.55, p=n.s.).
5.5.5. Control and Ownership
Participants indicated higher perceived control (treatment: M=5.00,Q1=4.50,Q3=6.00, baseline: M=5.00, Q1=3.00,Q3=6.00) and satisfaction with their control (treatment: M=5.00,Q1=2.75,Q3=6.00, baseline: M=5.00,Q1=3.00,Q3=5.25) under the treatment condition only if we look within-subjects (Figure 7a-b). Eleven participants had higher perceived control with Papers-to-Posts than the baseline tool (equal: 2, lower: 7), and 10 participants had higher satisfaction with Papers-to-Posts (equal: 4, lower: 6). On the other hand, participants did not demonstrate increased perceived ownership (treatment: M=6.00,Q1=5.75,Q3=6.00, baseline: M=6.00,Q1=5.00,Q3=6.00) or satisfaction with ownership (treatment: M=6,Q1=4.00,Q3=6.25, baseline: M=6.00,Q1=5.00,Q3=6.00) under the treatment condition (Figure 7c-d). Although they felt that they had more control over Papers-to-Posts, participants may not necessarily have felt that the LLM tool contributed more or less to the creation of the blog post.
5.5.6. Fulfillment of Design Goal 1
Looking at Figure 8a-b, we see that participants were more confident that all essential information was in the blog post (treatment: M=6.00,Q1=3.00,Q3=6.25, baseline: M=5.00,Q1=3.00,Q3=6.00) and found it easier to include essential information from the paper (treatment: M=6.00,Q1=5.00,Q3=6.00, baseline: M=5.00,Q1=4.75,Q3=6.00) using Papers-to-Posts. In addition, the interaction logs show that participants utilized Papers-to-Posts’s DG1-inspired affordance to select relevant paper bullet points and paragraphs to incorporate in the blog post. All participants adjusted the LLM’s pre-selected bullet points to some degree. On average, participants added or removed 28 bullet points and 13 paragraphs. We note that some added bullet points come from automatic selection of bullet points used to generate a new section rather than manual selection. Participants also provided an average of 1.70 custom bullet points.
In their short answers regarding aspects of Papers-to-Posts that they found difficult and helpful, 10 participants noted the paper outline being useful in some manner. Four participants described the outline as helpful for generating text. For instance, P15 noted, ”I could rewrite entire sections or direct it with just a few bullets.” Three participants mentioned that the outline was useful for reviewing the information in or structure of the paper. P4 commented, ”The outline was really helpful in providing a quick overview of what content I had written.” For two participants, the bullet points provided transparency around the origins of the initial LLM draft and generations. P6 explained, ”The checked-marked bullet points helped me understand where the pre-filled content on the left was from,” and P12 shared, ”I liked the bullet point features to see what was being fed in vs not.” Five participants also specified a difficulty related to the paper outline. Two participants were overwhelmed with the number of bullet points. P9 described the outline as ”really bloated and hard to navigate.” Three more participants were confused or encumbered by aspects of utilizing the bullet points. For example, P3 ”felt stressed by choosing between the entire paragraph at hand and the (presumably) AI-generated summary bullets,” and P20 did not realize that bullet points could come from different sections of the paper.
5.5.7. Fulfillment of Design Goal 2
Participants’ Likert responses regarding DG2 indicate that they found it easier to produce appropriate blog post text (treatment: M=6.00,Q1=5.00,Q3=6.00, baseline: M=5.00,Q1=4.00,Q3=6.00) and iterate on the blog post text (treatment: M=6.00,Q1=5.00,Q3=6.00, baseline: M=5.00,Q1=4.00,Q3=6.00) in the treatment condition (Figure 8c-d). Furthermore, the interaction logs for the treatment condition indicate that participants took advantage of DG2-based affordances for generating and modifying text for sub-tasks (Appendix 27). For an average of 6.85 generations per treatment condition, participants often utilized custom instructions (5.70 times), selected paragraphs or bullet points (6.35 times), and the context of prior blog post sections (6.00 times), though we note that these features were toggled on by default. In addition, they modified text 6.85 times on average, with the ”condense” (1.85 times), ”custom” (1.80 times), and ”simpler terms” (1.40 times) modifications being the most common.
In their short answers, 5 participants noted the usefulness of being able to provide custom instructions for either generating or modifying text using Papers-to-Posts. P9 described the custom instructions as providing a greater sense of control: ”…being able to include explicit instructions for the model to generate text from was helpful in being able to control the information in the text that it generated.” With respect to revising writing in particular, 12 participants noted in their short answers that Papers-to-Posts’s affordances for modifying text were useful. All of the preset modification buttons (”simpler terms”, ”condense”, ”expand”, ”more dramatic”, ”less dramatic”) were specifically mentioned as helpful affordances at least once. The most commonly noted one was the ”condense” button. For example, P12 commented, ”I sometimes put the modify text through multiple rounds of expand and contract, focusing on different parts. I found that helpful.” Two participants commented that they desired more preset modification buttons. P16 wanted a ”more technical” button, while P7 simply shared, ”I wish there are more default buttons for similar functions.” P7 then added, ”Or even let user to add their favorite rewrite button.” Interestingly, P2 found the preset modification buttons useful for understanding the LLM tool’s capabilities. They reflected, ”It was nice to know what options the tool already knew of, and would perhaps have success with. For example, I would probably not have gone to a blank-box tool and said, ’please make the following more dramatic’. I would have stuck to more basic things like ’please summarize the following’.”
Participants also discussed difficulties that they encountered in drafting and revising writing with Papers-to-Posts. Eight participants commented that they appreciated the simplicity of the baseline tool or that the baseline tool gave them more freedom over the instructions sent to the tool. For instance, P9 noted, ”the ’generate writing’ panel felt overwhelming in terms of features,” and P19 explained, ”It’s very open end so I felt like I got more freedom what to ask it to do.” Meanwhile, in the treatment condition, 4 participants expressed wanting to take specific actions that were not explicitly supported by Papers-to-Posts. As an example, 2 participants wanted assistance with checking for redundancies in the blog post. Another 2 wanted feedback on the blog post writing. Regarding the affordances for modifying text in particular, 8 participants noted a relative difficulty. A common issue, raised by 4 participants, was confusion about how or when to use certain affordances. As an example, P3 ”didn’t really understand if the custom instructions only applied for the ’custom modification’ option,” when it in fact would be applied to any modification action. Meanwhile, P6 was unsure of ”when to use which one of the six [modification] options.” Another recurrent issue noted by 5 participants was a modification affordance not working as desired. For P20, setting the desired length of the modified writing to one paragraph still resulted in three, as the back-end prompts are not guaranteed to work as planned. For P12, modifications were sometimes too exaggerated, while for P16, a custom modification to make writing more technical was too subdued. Despite these comments about feeling restricted or confused by Papers-to-Posts, participants still reported higher perceived control and satisfaction with control in the treatment condition, which we discuss in Section 5.5.5.
5.5.8. Extent of Change to Blog Post
Interestingly, we found that participants made larger changes to the initial LLM draft blog post when using Papers-to-Posts as compared to the baseline tool. In Figure 9a, for each condition, we show the Levenshtein distance, a common metric for measuring the amount of editing done to text, between participants’ initial LLM draft blog posts and their blog posts at the point at which the writer would feel comfortable publishing it publicly. For exploratory statistical testing, we employed a paired-samples t-test and observed that this Levenshtein distance is significantly higher with Papers-to-Posts (M=3977.05,SD=1146.46) compared to the baseline (M=2764.15,SD=1042.16) (Paired-Samples t-Test, t(19)=4.31, p¡.0005).
In Figure 9a, we show how the Levenshtein distance progressed over time, with the removal of one outlier. The difference in Levenshtein distance does not appear attributable to the time spent on the task or the change in length of the blog post. Participants did not spend substantially more time to reach the point at which they were comfortable publicly publishing the blog post using Papers-to-Posts (M=49.37, Q1=35.43, Q3=59.54) as compared to the baseline (M=46.81, Q1=33.92,Q3=56.60) (Appendix 25). Furthermore, we saw that participants changed the length of their blog posts, measured in characters, less using Papers-to-Posts (M=-162.00, Q1=-1126.25, Q3=297.00) than using the baseline (M=1971.00, Q1=965.25,Q3=2514.25) (Appendix 25(b)). Thus, we suspected that Papers-to-Posts may have required a decreased number of writing actions (inserting, deleting, replacing text in text fields) in order to achieve a given increase in Levenshtein distance. We investigated this potential reduction in effort in the follow-up deployment study.
6. Deployment Study
Through the controlled lab study, we observed that participants produced final blog posts that were significantly more distant from their initial LLM-generated drafts when using Papers-to-Posts than the baseline tool. We wanted to further understand how participants in a more realistic setting perceive Papers-to-Posts versus the baseline tool (RQ2), and whether Papers-to-Posts would lead participants to change their provided drafts more in a given time period (RQ1). For this, we conducted a deployment study where participants wrote blog posts on their own time with the explicit goal of posting the article publicly and sharing it with others.
6.1. Participants
We recruited 26 participants (M: 17, W: 8, undisclosed: 1) through academic social networks and institutional mailing lists. They were compensated with $50 over PayPal for each blog post that they wrote, with a maximum of two blog posts. Participants were predominantly early career researchers (undergraduate student: 1, master’s student: 4, PhD student: 13, postdoc: 2, professor: 2, industry researcher: 4). All participants conducted research in an area of computer science, except for one who studied cognitive science. The majority of the participants had interacted with an LLM several times (¿10 sittings: 21, 6-10 sittings: 4, 2-5 sittings: 1). The participants had varied experience in writing blog posts (0 posts: 16, 1-2 posts: 9, 3-10 posts: 1). They were required to have authored at least one research paper (1-2 papers: 6, 3-10 papers: 11, ¿10 research papers: 9) and to be more interested than not (at least 5 on a 7-point Likert scale where 7 indicates the highest interest) in writing research-paper blog posts. We also required participants to be willing to share their final blog posts with others if they were able to reach the point at which they would feel comfortable doing so. Lastly, the papers that the participants want to write about had to be compatible with our tools’ PDF parser.
6.2. Study Conditions
While the treatment and baseline conditions were largely the same as in the lab study, several minor usability issues identified in the lab study were addressed before proceeding with the deployment study. For example, there were better instructions and less intrusive highlight colors (Appendix 29). We also replaced GPT-4 and GPT-4-32k with the newer GPT-4-0125-preview model with a larger context window that became available after the lab study. This enabled Papers-to-Poststo support longer inputs. In the deployment study, participants could freely close and reopen the tool in their browsers to work on their drafts when convenient. Participants could save their progress across sessions with a newly added save button. We note that, due to the manner in which text was logged in the deployment study, the tool in both conditions had a lag if participants pasted a chunk of text more then a couple paragraphs long. Lastly, the treatment tool now automatically parsed the input paper.
The majority of the 37 papers about which participants wrote were published in 2023 or 2024 (26 papers). The two oldest papers were from 2019 and 2008. The average parsed paper length across conditions differed by less than 5% (baseline: 37064 characters, treatment: 35845 characters).
6.3. Study Procedure
The task for the deployment study was to share a blog post with others, which can require significantly higher effort than the controlled lab study, which only required voicing comfort in publicly posting the blog post. To control for the study length and ease of recruitment, we opted for a between-subjects study so that each participant only had to write one blog post. At the same time, we allowed participants who wished to participate in the study twice (N=11/26 participants) to use both tools each time in randomized, counter-balanced order. When analyzing the results we treated them as non-paired data. Eighteen participants completed the baseline condition, and 19 participants completed the treatment condition. For participants who wrote two blog posts, their selected papers were assigned randomly to each condition. Participants had up to six days to complete a research-paper blog post using the assigned tool on their own time. We provided participants with a video tutorial on the assigned tool, a link to the paper about which they would be writing, and a link to the assigned tool. We noted that their aim should be to share the final blog post with others, but if they felt that they could not reach the point of being comfortable with sharing it with others, they could let us know and still be compensated. Participants were required to spend at least 20 minutes using the tool to write the blog post. To better understand how participants utilized the two systems, we recorded their interactions, including prompting, inserting, deleting and replacing text. In addition, we also took periodic snapshots of the drafts whenever they (re)opened the systems in their browsers and every 60 seconds during active usage.
Once participants were satisfied with their draft, participants shared a blog post based on their drafts. We provided instructions on how to create blog posts on the Medium platform,222https://medium.com/ but they were allowed to use other publishing platforms if they preferred. They then emailed us the blog post link and a screenshot or link showcasing evidence that they shared the blog post with others (e.g., tweet, email to co-authors, slack channel message, LinkedIn post). Only 2 of the 26 participants (one under both conditions and one under the treatment condition) alerted us to the fact that they could not reach the point at which they felt comfortable sharing the blog post with others in the allotted time.333We note that some participants indicated that the blog post was part of an experiment when sharing with others, which may have increased their willingness to share it; regardless, they were willing to share a post at the end of the study. Participants then completed the same survey from the lab study regarding their experience with the tool. We note that four participants who we realized had not completed 20 minutes of interaction were asked to do so after completing the survey. The only difference was the removal of a few questions irrelevant to this study: a question regarding how much additional time would be needed to feel comfortable posting the blog post publicly and questions regarding a direct comparison between the two tools. Lastly, for the 15 participants who were willing, the first author engaged them in a 15-minute semi-structured interview to discuss their experience with the tool.
6.4. Results
We did not plan to investigate statistical significance in this between-subjects study, given power analyses indicating that we would need a larger sample size than we could obtain to see a significant result for our metric of Levenshtein distance (power=0.95, alpha=0.05, effect size=0.45; requires 29 participants per condition) or blog post satisfaction (power=0.95, alpha=0.05, effect size=0.98; requires 83 participants per condition). Throughout our analysis of the deployment study logs, we exclude the three study runs in which the participant made no change to the initial draft, as there is no interaction in producing the blog post. We also exclude six study runs in which the participant claimed to have made edits outside the tool beyond small things like grammar, formatting, and adding images; this is to avoid drawing conclusions based on missing data about effort spent. The resulting log data used thus includes 28 study runs (treatment: 14, baseline: 14). The first author conducted inductive thematic analysis (Braun and Clarke, 2006) of the participants’ survey and interview comments.
6.4.1. RQ1: Blog Post Iteration
When given a time constraint in the lab study, we observed that participants made larger changes to the tool’s initial draft with Papers-to-Posts than with the baseline tool. The deployment study allowed us to investigate if this evident decrease in effort could be reflected in a realistic setting and in terms of other metrics of effort (i.e., number of writing actions). We found that participants using Papers-to-Posts took fewer writing actions in order to create one unit of change in Levenshtein distance to the blog post (Figure 9(c)). We also observed, in alignment with the lab study, that Papers-to-Posts led to greater changes in Levenshtein distance for a given amount of active time, which we define as time taking any action in the tool (Figure 9(b)). Therefore, when using Papers-to-Posts as compared to the baseline, participants experienced reduced effort in iterating on the blog post with respect to both writing actions needed and time needed.
Once more, this is despite the fact that participants did not spend substantially more active time using Papers-to-Posts (M=22.00, Q1=17.25,Q3=34.00) than the baseline (M=22.50,Q1=16.00,Q3=27.00) (Appendix 26(a)). Furthermore, we again saw that participants changed the length of their blog posts, measured in characters, less using Papers-to-Posts (M=-5.5, Q1=-525.25, Q3=961.00) than using the baseline (M=710.00, Q1=47.25, Q3=4261.25) (Appendix 26(b)). In fact, Figure 9 shows that the maximum active time and writing actions taken by a baseline participant is higher than that of a treatment participant. This provides some evidence that, for writers who tend to spend more time than average on the task, Papers-to-Posts may help them to reduce the time they spend on it.
Participants’ survey and interview comments reflected this discrepancy between the two tools. With respect to the baseline, participants most commonly cited the initial draft as a helpful aspect of the tool. Nine participants under the baseline condition specifically noted that it was useful as a starting point. Focusing on the initial draft, baseline participants often spoke of editing or adding to it rather than replacing it. For example, P20 explained, ”It’s easier to critique than it is to generate,” and P26 shared, ”It was very useful to have a straw-person of what the blog post could look like (generated by the LLM). I mostly used the draft as a place to edit and steer in the direction I wanted the post to take.”
On the other hand, 15 participants noted the selectable paper bullet points and paragraphs as a helpful feature of Papers-to-Posts. Focusing on the feature designed to generate full blog post sections, participants using Papers-to-Posts often reported doing just that. As an example, P12 commented that the ”outline [to] text generation process is very helpful to get the first draft of the post.” Furthermore, because the tool shows users what bullet points the tool used to generate the initial draft, participants may have been inclined to inspect and iterate upon this selection. This is illustrated by P20, who observed, ”The automatic selection of bullet points to include did not capture most of the information I wanted it to, and so required a lot of manual selection of bullet points, but once I selected those, the generation made it easy to write.” Participants also frequently mentioned the utility of the Papers-to-Posts’s modification abilities. P18 reflected, ”The ”MODIFY WRITING” section is helpful. I like that the interface contains this particular section to support iterating on existing text.” By facilitating modifications to the initial draft, Papers-to-Posts may have encouraged participants to iterate more heavily on the initial draft.
6.4.2. RQ2: Real-World Tool Perceptions
We evaluated participants’ survey results in the same manner as in the lab study (see Section 5.5). With respect to satisfaction with the final blog post quality, we observed no difference in the median satisfaction under Papers-to-Posts(M=6.00,Q1=5.00,Q3=6.00) versus the baseline (M=6.00,Q1=5.25, Q3=6.00) (Figure 5j-m). This discrepancy with the lab study may be related to the fact that participants were not writing under a time constraint in the deployment study. With sufficient time to write the blog post to their liking (only 3 blog posts were not shared with others), participants may not have noticed a pronounced difference in blog post quality between conditions.
Similarly to the lab study, we observed slightly greater satisfaction with the tool under the treatment condition (M=5.60,Q1=5.00, Q3=6.40) than the baseline condition (M=5.40,Q1=4.05,Q3=5.80) (Figure 5n-r). In alignment, four participants who participated twice expressed a preference for Papers-to-Posts, whereas no one shared a preference for the baseline tool. P19 remarked, ”I mean, honestly, I like the first tool better than the second tool… the second tool is a bit simplistic.” However, the complexity and larger associated learning curve of Papers-to-Posts may have decreased its chances of being widely preferred over the baseline. For example, P15 commented, ”It was a lot of information on the interface to deal with. I would prefer less moving things in front of me.” Nevertheless, we once again found slightly lower cognitive load under the treatment condition (M=15.00,Q1=9.50,Q3=17.50) in comparison to the baseline condition (M=15.50,Q1=10.75,Q3=18.75) (Figure 6e-h). This may be related to the fact that, while Papers-to-Posts required understanding a more complex interface, the baseline tool placed greater demands on participants in terms of writing. P24 summarized, ”In this [baseline] version, you have greater autonomy regarding what needs to be done. While this is beneficial, it also requires more effort and planning.”
As shown in Figure 7e-h, participants indicated higher satisfaction with control (treatment: M=5.00,Q1=3.00,Q3=6.00, baseline: M=3.00,Q1=3.00,Q3=4.75) in the treatment condition like in the lab study, but they showed no difference in median satisfaction with respect to perceived control (treatment: M=5.00,Q1=5.00,Q3=6.00, baseline: M=5.00,Q1=3.00,Q3=5.00), perceived ownership (treatment: M=6.00,Q1=4.50,Q3=6.00, baseline: M=6.00,Q1=6.00,Q3=6.00) or satisfaction with ownership (treatment: M=6.00,Q1=4.00,Q3=6.00, baseline: M=6.00,Q1=5.00,Q3=6.00) in the treatment condition. Participants may have been more satisfied with their control of the tool without perceiving a higher control using Papers-to-Posts in the sense that they were comfortable with letting the tool have a certain amount of control in exchange for less effort to generate text.
Overall, we observed benefits from Papers-to-Posts in fulfilling Papers-to-Posts’s design goals. Regarding the first design goal (DG1), participants indicated only a slight improvement in feeling confident that all essential information was covered with Papers-to-Posts (M=5.00,Q1=4.50,Q3=6.00) as opposed to the baseline (M=5.00,Q1=4.25,Q3=6.00). Given that the automatic method for parsing the paper PDF in the deployment study was likely to make more errors than the partially Wizard-of-Oz method for parsing papers in the lab study, it makes sense that participants would now see a reduced benefit from Papers-to-Posts’s paper outline in terms of ensuring all relevant information is included. Multiple participants under both conditions noted that the parsing of their papers was incomplete or flawed. However, participants still indicated an increased ease in incorporating information from the paper in the blog post with Papers-to-Posts (M=6.00,Q1=5.00,Q3=6.00) compared to the baseline (M=4.50,Q1=3.25,Q3=5.75). For instance, P14 commented, ”I liked being able to check the boxes of the portions of the paper I wanted to include in each section.”
With respect to our second design goal (DG2), there was no difference in the median for how well Papers-to-Posts (M=5.00,Q1=4.00, Q3=6.00) versus the baseline (M=5.00,Q1=5.00,Q3=6.00) facilitated production of appropriate text, unlike in the lab study (Figure 8e-h). Nevertheless, in their qualitative feedback, several participants noted how Papers-to-Posts’s generation and modification features helped them to create more appropriate text. For instance, P24 explained that the ”preset custom instructions was very useful because can need to be very specific using these tools,” and P14 noted that the ”provided buttons for making text more or less dramatic was nice.” Furthermore, in their survey responses, participants indicated that Papers-to-Posts (M=6.00,Q1=4.50,Q3=6.00) facilitated iterating on the blog post more so than the baseline (M=5.00,Q1=3.50,Q3=6.00), which is reflected in the lower number of writing actions participants took for a unit change in their initial drafts that we observed in Figure 9(c).
7. Discussion and Design Implications
Based on the results of the lab and deployment studies, we reflect on design implications for other human-LLM scientific long-document summarization tools with long-form outputs.
7.1. Post-Editing is not Enough
Both Papers-to-Posts and the baseline tool included initial LLM-generated drafts generated with powerful LLMs, GPT-4 and GPT-4-32k respectively. Thus, the existence of an initial LLM draft cannot be the primary reason why participants preferred the blog post generated with Papers-to-Posts as compared to the baseline tool. While the participants had the option to simply post-edit the initial draft and refrain from interacting further with the LLM, only two baseline and two treatment participants across the lab and deployment studies did so. This suggests that additional user interfaces and interaction designs can provide better scaffolding supports to further improve upon LLM-generated text compared to simple post-editing.
7.2. Long-Document Outlines
While blog post writing has additional goals and constraints compared to general document summarization, our work also has implications that may be useful for the task of general human-AI long document summarization. Specifically, a main challenge here is the high costs of human summarizer reviewing, locating, and utilizing different essential information in a long document. Prior work in human-LLM summarization has attempted to address this challenge in various manners, including presenting visualizations for the document to be summarized (Zhang et al., 2023b, 2011), dividing the document into chunks (Cheng et al., 2023), and providing document highlights for information already included in the summary (Norkute et al., 2021). However, to the best of our knowledge, no prior work has explored the use of generated outlines to help users review the document to be summarized.
In this work, participants found it easier to find and incorporate essential information from the document using Papers-to-Posts and participants consistently noted the selectable bullet points as one of the most useful features. We found that, given this tool, participants were able to make more change to their initial drafts for a given amount of time or writing actions. We therefore note that presenting LLM-generated outlines of scientific long documents may make the planning step of writing their summaries easier. As participants faced confusion about the connection between the outline and the paper text, we also recommend working to make this connection as clear as possible. How detailed the outline should be remains uncertain. While several participants appreciated the bullet points, a couple mentioned that the outline was overwhelming. As one participant proposed, a hierarchy of bullet points may be more useful.
7.3. Transparent Affordances for Long-form Writing Sub-Tasks
In contrast with shorter forms of scientific writing such as tweetorials (Long et al., 2023) and extended metaphors (Kim et al., 2023) explored in prior work on human-LLM scientific writing, longer scientific texts such as research-paper blog posts require completing a number of sub-tasks common to long-form science communication. For example, because research-paper blog posts are often for broader audiences than the paper itself, Papers-to-Posts includes a ”simpler terms” option to make large swaths of text more digestible for such readers. As another example, because different sections of a research-paper blog post may serve different purposes, we pre-populated the custom instructions for each of the initial draft sections (e.g., for the introduction section, we ask the tool to present a hook, a high-level description of the problem and why it is interesting, etc.). These transparent, section-specific instructions could then be verified or adjusted by the user. Several participants commented on the usefulness of Papers-to-Posts’s different affordances for streamlining different generation and modification sub-tasks such as these. Thus, we observe that similar tools may assist users by including affordances for sub-tasks related to generating and modifying long-form output.
8. Limitations and Future Work
The majority of the participants were PhD students in computer science. Although many of the takeaways discussed in this paper may generalize to other groups of researchers, it would be interesting to see how more diverse populations surface unique needs from the tool. For example, researchers in the field of computer science tend to be better versed in artificial intelligence than those in other fields, so they may not need as much guidance in utilizing an LLM writing tool. Different areas of research may also have very different traditional paper and blog post structures compared to computer science, which may benefit from different tool affordances. Furthermore, PhD students are prime candidates for writing blog posts, which can help them raise awareness of their work as they are starting their research careers. However, more senior researchers may desire specific features for writing their blog posts, such as a feature to coordinate discussion of multiple works or a feature to connect an older work to the current research landscape. We also leave it to future work to investigate how different aspects of the input paper such as length, recency, and author position may affect the support that users want from the tool.
9. Ethical Considerations
There are a few ethical considerations that we would like to note regarding this work. For one, it is well-known that LLMs may hallucinate information included in their generated text (Bubeck et al., 2023). We mitigate this risk by focusing on researchers summarizing their own work, which they should know well; for example, multiple participants noted inaccuracies or exaggerations in generated text, but most were able to correct those issues. A couple participants raised concerns that they had not caught all the errors by the end of their session, but as one of them noted, with more time they could be thorough in ensuring that the text is correct. We believe that LLM usage will increase in the future, and it is important to study effective design. Another ethical concern is whether using an LLM tool to write research-paper blog posts could lead to the loss of important details from the paper. P3 in the baseline condition raised this point: ”…it seems like the easier you make it, i.e., the better the initial summary, the less likely I am to myself remember those things that happen to be missing.” That said, participants overall felt more confident that all essential information was covered in their blog posts with Papers-to-Posts, so this work provided one step in the direction of support for reducing missing information. Furthermore, an LLM tool for writing research-paper blog posts may contribute to making science more accessible in society. This may raise concerns of reducing paper readership. However, blog posts do not provide the same amount of information as papers and do not go through peer review, research papers still hold their own value, and blog posts can act as catalysts for paper reading. Finally, researchers may not acknowledge the use of AI in writing their blog posts, which could mislead readers, particularly if the researcher has not verified the blog post contents. For the participants who shared their blog posts, we found that many chose to disclose its use, but to ensure better transparency around its use, future versions of Papers-to-Posts may highlight the provenance of the output writing (e.g., by visualizing the writer’s interaction with the LLM-powered tool (Naimul Hoque et al., 2023)).
10. Conclusion
We introduce Papers-to-Posts, an LLM-powered research-paper blog post writing tool, which effectively helps researchers plan, draft, and revise their blog posts, Through a formative study (N=6) with an LLM technology probe for this task, we identified the following challenges faced by writers: they 1) appreciated a paper outline for finding and selecting relevant information from the paper but faced difficulties in navigating it and 2) tried to provide recurrent instructions for different kinds of sub-tasks to the LLM without specific affordances to support these sub-tasks. We formulated design goals to address each of these challenges, namely to help writers 1) plan their writing by facilitating review of the long document’s contents and selection of essential information to include in the writing and 2) draft and revise their writing by streamlining commonly desired instructions to the LLM for generating and modifying text. We validated that our LLM-powered tool Papers-to-Posts addresses these goals and helps authors write blog posts for their own papers through two user studies comparing Papers-to-Posts against a baseline providing an LLM blog post draft and free-form LLM prompting. In a within-subjects, time-controlled lab study (N=20 participants), we found that Papers-to-Posts significantly increased writer satisfaction with the output blog post, without increasing task completion time or cognitive load. We also observe that Papers-to-Posts led to significantly greater change to the initial LLM draft blog post, indicating that this increased satisfaction may be due to the tool facilitating iteration on the blog post. A field deployment study (N=37 blog posts, 26 participants) showed that in a more realistic setting Papers-to-Posts again resulted in greater change in the blog post for a given amount of editing time or number of editing actions. We hope that this work may inform the design of future human-LLM tools for scientific long-document summarization into long-form output.
Acknowledgements.
This research was supported by the Allen Institute for Artificial Intelligence (AI2). The authors thank the Semantic Scholar team for helpful feedback on this work and the participants who made this work possible.References
- (1)
- August et al. (2020) Tal August, Lauren Kim, Katharina Reinecke, and Noah A Smith. 2020. Writing strategies for science communication: Data and computational analysis. In Proceedings of the 2020 Conference on Empirical Methods in Natural Language Processing (EMNLP). 5327–5344.
- Avinesh et al. (2018) PVS Avinesh, Carsten Binnig, Benjamin Hättasch, Christian M Meyer, and Orkan Özyurt. 2018. Sherlock: A System for Interactive Summarization of Large Text Collections. Proc. VLDB Endow. 11, 12 (2018), 1902–1905.
- Avinesh and Meyer (2017) PVS Avinesh and Christian M Meyer. 2017. Joint optimization of user-desired content in multi-document summaries by learning from user feedback. In Proceedings of the 55th Annual Meeting of the Association for Computational Linguistics (Volume 1: Long Papers). 1353–1363.
- Bayatmakou et al. (2022) Farnoush Bayatmakou, Azadeh Mohebi, and Abbas Ahmadi. 2022. An interactive query-based approach for summarizing scientific documents. Information Discovery and Delivery 50, 2 (2022), 176–191.
- Blanchard (2011) Antoine Blanchard. 2011. 3.4 Science Blogs in Research and Popularization of Science: Why, how and for whom? Common knowledge: The challenge of transdisciplinarity (2011), 219–232.
- Bohn and Ling (2021) Tanner Bohn and Charles X Ling. 2021. Hone as you read: A practical type of interactive summarization. arXiv preprint arXiv:2105.02923 (2021).
- Braun and Clarke (2006) Virginia Braun and Victoria Clarke. 2006. Using thematic analysis in psychology. Qualitative research in psychology 3, 2 (2006), 77–101.
- Bubeck et al. (2023) Sébastien Bubeck, Varun Chandrasekaran, Ronen Eldan, Johannes Gehrke, Eric Horvitz, Ece Kamar, Peter Lee, Yin Tat Lee, Yuanzhi Li, Scott Lundberg, et al. 2023. Sparks of artificial general intelligence: Early experiments with gpt-4. arXiv preprint arXiv:2303.12712 (2023).
- Burns et al. (2003) Terry W Burns, D John O’Connor, and Susan M Stocklmayer. 2003. Science communication: a contemporary definition. Public understanding of science 12, 2 (2003), 183–202.
- Cai et al. (2022) Pengshan Cai, Fei Liu, Adarsha Bajracharya, Joe Sills, Alok Kapoor, Weisong Liu, Dan Berlowitz, David Levy, Richeek Pradhan, and Hong Yu. 2022. Generation of patient after-visit summaries to support physicians. In Proceedings of the 29th International Conference on Computational Linguistics (COLING).
- Cheng et al. (2023) Chen Cheng, Ziang Li, Zhenhui Peng, and Quan Li. 2023. ALens: An Adaptive Domain-Oriented Abstract Writing Training Tool for Novice Researchers. arXiv preprint arXiv:2308.03989 (2023).
- Cheng et al. (2022) Ruijia Cheng, Alison Smith-Renner, Ke Zhang, Joel R Tetreault, and Alejandro Jaimes. 2022. Mapping the design space of human-ai interaction in text summarization. arXiv preprint arXiv:2206.14863 (2022).
- Chung et al. (2022) John Joon Young Chung, Wooseok Kim, Kang Min Yoo, Hwaran Lee, Eytan Adar, and Minsuk Chang. 2022. TaleBrush: Sketching Stories with Generative Pretrained Language Models. Proceedings of the 2022 CHI Conference on Human Factors in Computing Systems (2022). https://api.semanticscholar.org/CorpusID:247625751
- CMU ([n. d.]) CMU. [n. d.]. CMU. https://blog.ml.cmu.edu/submissions/.
- Coenen et al. (2021) Andy Coenen, Luke Davis, Daphne Ippolito, Emily Reif, and Ann Yuan. 2021. Wordcraft: a human-ai collaborative editor for story writing. arXiv preprint arXiv:2107.07430 (2021).
- Dang et al. (2022) Hai Dang, Karim Benharrak, Florian Lehmann, and Daniel Buschek. 2022. Beyond text generation: Supporting writers with continuous automatic text summaries. In Proceedings of the 35th Annual ACM Symposium on User Interface Software and Technology. 1–13.
- Ding et al. (2023) Zijian Ding, Alison Smith-Renner, Wenjuan Zhang, Joel R Tetreault, and Alejandro Jaimes. 2023. Harnessing the power of LLMs: Evaluating human-AI text co-creation through the lens of news headline generation. arXiv preprint arXiv:2310.10706 (2023).
- El-Kassas et al. (2021) Wafaa S El-Kassas, Cherif R Salama, Ahmed A Rafea, and Hoda K Mohamed. 2021. Automatic text summarization: A comprehensive survey. Expert systems with applications 165 (2021), 113679.
- Flower and Hayes (1981) Linda S. Flower and J. R. Hayes. 1981. A Cognitive Process Theory of Writing. College Composition and Communication 32 (1981), 365–387. https://api.semanticscholar.org/CorpusID:18484126
- Gao et al. (2020) Yang Gao, Christian M Meyer, and Iryna Gurevych. 2020. Preference-based interactive multi-document summarisation. Information Retrieval Journal 23, 6 (2020), 555–585.
- Gehrmann et al. (2019) Sebastian Gehrmann, Hendrik Strobelt, Robert Krüger, Hanspeter Pfister, and Alexander M Rush. 2019. Visual interaction with deep learning models through collaborative semantic inference. IEEE transactions on visualization and computer graphics 26, 1 (2019), 884–894.
- Gero et al. (2021) K. Gero, Vivian Liu, Sarah Huang, Jennifer Lee, and Lydia B. Chilton. 2021. What Makes Tweetorials Tick: How Experts Communicate Complex Topics on Twitter. Proceedings of the ACM on Human-Computer Interaction 5 (2021), 1 – 26. https://api.semanticscholar.org/CorpusID:239020743
- Gero et al. (2022) Katy Ilonka Gero, Vivian Liu, and Lydia Chilton. 2022. Sparks: Inspiration for science writing using language models. In Designing interactive systems conference. 1002–1019.
- Ghazvininejad et al. (2017) Marjan Ghazvininejad, Xing Shi, Jay Priyadarshi, and Kevin Knight. 2017. Hafez: an interactive poetry generation system. In Proceedings of ACL 2017, System Demonstrations. 43–48.
- Goldfarb-Tarrant et al. (2019) Seraphina Goldfarb-Tarrant, Haining Feng, and Nanyun Peng. 2019. Plan, write, and revise: an interactive system for open-domain story generation. arXiv preprint arXiv:1904.02357 (2019).
- Google ([n. d.]) Google. [n. d.]. BARD. https://bard.google.com/chat.
- Hart and Staveland (1988) Sandra G Hart and Lowell E Staveland. 1988. Development of NASA-TLX (Task Load Index): Results of empirical and theoretical research. In Advances in psychology. Vol. 52. Elsevier, 139–183.
- Holmberg and Thelwall (2014) Kim Holmberg and Mike Thelwall. 2014. Disciplinary differences in Twitter scholarly communication. Scientometrics 101 (2014), 1027–1042.
- Jarreau (2015) Paige Brown Jarreau. 2015. All the science that is fit to blog: An analysis of science blogging practices. Louisiana State University and Agricultural & Mechanical College.
- Jucan and Jucan (2014) Mihaela Sabina Jucan and Cornel Nicolae Jucan. 2014. The power of science communication. Procedia-Social and Behavioral Sciences 149 (2014), 461–466.
- Kim et al. (2023) Jeong Chul Kim, Sang Guk Suh, Lydia B. Chilton, and Haijun Xia. 2023. Metaphorian: Leveraging Large Language Models to Support Extended Metaphor Creation for Science Writing. Proceedings of the 2023 ACM Designing Interactive Systems Conference (2023). https://api.semanticscholar.org/CorpusID:259376400
- Kocielnik et al. (2019) Rafal Kocielnik, Saleema Amershi, and Paul N Bennett. 2019. Will you accept an imperfect ai? exploring designs for adjusting end-user expectations of ai systems. In Proceedings of the 2019 CHI Conference on Human Factors in Computing Systems. 1–14.
- Koh et al. (2022) Huan Yee Koh, Jiaxin Ju, Ming Liu, and Shirui Pan. 2022. An empirical survey on long document summarization: Datasets, models, and metrics. ACM computing surveys 55, 8 (2022), 1–35.
- Kudos ([n. d.]) Kudos. [n. d.]. Kudos. https://info.growkudos.com/landing/researchers-2023.
- Laban et al. (2023) Philippe Laban, Jesse Vig, Marti A Hearst, Caiming Xiong, and Chien-Sheng Wu. 2023. Beyond the Chat: Executable and Verifiable Text-Editing with LLMs. arXiv preprint arXiv:2309.15337 (2023).
- Lai et al. (2022) Vivian Lai, Alison Smith-Renner, Ke Zhang, Ruijia Cheng, Wenjuan Zhang, Joel Tetreault, and Alejandro Jaimes. 2022. An exploration of post-editing effectiveness in text summarization. arXiv preprint arXiv:2206.06383 (2022).
- Long et al. (2023) Tao Long, Dorothy Zhang, Grace Li, Batool Taraif, Samia Menon, Kynnedy Simone Smith, Sitong Wang, Katy Ilonka Gero, and Lydia B. Chilton. 2023. Tweetorial Hooks: Generative AI Tools to Motivate Science on Social Media. ArXiv abs/2305.12265 (2023). https://api.semanticscholar.org/CorpusID:258833277
- Lu et al. (2023) Guang Lu, Sylvia B Larcher, and Tu Tran. 2023. Hybrid Long Document Summarization using C2F-FAR and ChatGPT: A Practical Study. arXiv preprint arXiv:2306.01169 (2023).
- Luc et al. (2021) Jessica GY Luc, Michael A Archer, Rakesh C Arora, Edward M Bender, Arie Blitz, David T Cooke, Tamara Ni Hlci, Biniam Kidane, Maral Ouzounian, Thomas K Varghese Jr, et al. 2021. Does tweeting improve citations? One-year results from the TSSMN prospective randomized trial. The Annals of thoracic surgery 111, 1 (2021), 296–300.
- Mahrt and Puschmann (2014) Merja Mahrt and Cornelius Puschmann. 2014. Science blogging: An exploratory study of motives, styles, and audience reactions. Journal of science Communication 13, 03 (2014), A05.
- Medium ([n. d.]) Medium. [n. d.]. Medium. https://blog.medium.com/best-practices-for-writing-on-medium-386506ae62b9.
- Mehlenbacher (2019) Ashley Mehlenbacher. 2019. Science communication online: Engaging experts and publics on the internet. The Ohio State University Press.
- Mirowski et al. (2023) Piotr Mirowski, Kory W Mathewson, Jaylen Pittman, and Richard Evans. 2023. Co-Writing Screenplays and Theatre Scripts with Language Models: Evaluation by Industry Professionals. In Proceedings of the 2023 CHI Conference on Human Factors in Computing Systems. 1–34.
- Moramarco et al. (2021) Francesco Moramarco, Alex Papadopoulos Korfiatis, Aleksandar Savkov, and Ehud Reiter. 2021. A preliminary study on evaluating consultation notes with post-editing. arXiv preprint arXiv:2104.04402 (2021).
- Naimul Hoque et al. (2023) Md Naimul Hoque, Tasfia Mashiat, Bhavya Ghai, Cecilia Shelton, Fanny Chevalier, Kari Kraus, and Niklas Elmqvist. 2023. The HaLLMark Effect: Supporting Provenance and Transparent Use of Large Language Models in Writing through Interactive Visualization. arXiv e-prints (2023), arXiv–2311.
- Norkute et al. (2021) Milda Norkute, Nadja Herger, Leszek Michalak, Andrew Mulder, and Sally Gao. 2021. Towards explainable AI: Assessing the usefulness and impact of added explainability features in legal document summarization. In Extended Abstracts of the 2021 CHI Conference on Human Factors in Computing Systems. 1–7.
- NotionAI ([n. d.]) NotionAI. [n. d.]. NotionAI. https://www.notion.so/product.
- OpenAI ([n. d.]) OpenAI. [n. d.]. chatGPT. https://chat.openai.com/.
- Patrick Dunleavy ([n. d.]) Patrick Dunleavy. [n. d.]. LSE. https://blogs.lse.ac.uk/impactofsocialsciences/2016/01/25/how-to-write-a-blogpost-from-your-journal-article/.
- Singh et al. (2022) Nikhil Singh, Guillermo Bernal, Daria Savchenko, and Elena L. Glassman. 2022. A Selective Summary of Where to Hide a Stolen Elephant: Leaps in Creative Writing with Multimodal Machine Intelligence. ACM Transactions on Computer-Human Interaction (2022). https://api.semanticscholar.org/CorpusID:247085270
- UWaterloo ([n. d.]) UWaterloo. [n. d.]. UWaterloo. https://uwaterloo.ca/writing-and-communication-centre/writing-academic-blog-posts.
- Van Someren et al. (1994) Maarten Van Someren, Yvonne F Barnard, and J Sandberg. 1994. The think aloud method: a practical approach to modelling cognitive. London: AcademicPress 11, 6 (1994).
- Wang et al. (2023a) Sitong Wang, Samia Menon, Tao Long, Keren Henderson, Dingzeyu Li, Kevin Crowston, Mark Hansen, Jeffrey V Nickerson, and Lydia B Chilton. 2023a. Reelframer: Co-creating news reels on social media with generative ai. arXiv preprint arXiv:2304.09653 (2023).
- Wang et al. (2023b) Sitong Wang, Zheng Ning, Anh Truong, Mira Dontcheva, Dingzeyu Li, and Lydia B Chilton. 2023b. PodReels: Human-AI Co-Creation of Video Podcast Teasers. arXiv preprint arXiv:2311.05867 (2023).
- Williams et al. (2022) Spencer Williams, Ridley Jones, Katharina Reinecke, and Gary Hsieh. 2022. An HCI Research Agenda for Online Science Communication. Proceedings of the ACM on Human-Computer Interaction 6 (2022), 1 – 22. https://api.semanticscholar.org/CorpusID:253459983
- Xie et al. (2023) Yujia Xie, Xun Wang, Si-Qing Chen, Wayne Xiong, and Pengcheng He. 2023. Interactive Editing for Text Summarization. arXiv preprint arXiv:2306.03067 (2023).
- Yan et al. (2011) Rui Yan, Jian-Yun Nie, and Xiaoming Li. 2011. Summarize what you are interested in: An optimization framework for interactive personalized summarization. In Proceedings of the 2011 conference on empirical methods in natural language processing. 1342–1351.
- Zarinbal et al. (2019) Marzieh Zarinbal, Azadeh Mohebi, Hesamoddin Mosalli, Razieh Haratinik, Zahra Jabalameli, and Farnoush Bayatmakou. 2019. A new social robot for interactive query-based summarization: Scientific document summarization. In Interactive Collaborative Robotics: 4th International Conference, ICR 2019, Istanbul, Turkey, August 20–25, 2019, Proceedings 4. Springer, 330–340.
- Zhang et al. (2023b) Xiaoyu Zhang, Jianping Li, Po-Wei Chi, Senthil Chandrasegaran, and Kwan-Liu Ma. 2023b. ConceptEVA: Concept-Based Interactive Exploration and Customization of Document Summaries. In Proceedings of the 2023 CHI Conference on Human Factors in Computing Systems. 1–16.
- Zhang et al. (2011) Yi Zhang, Dingding Wang, and Tao Li. 2011. iDVS: an interactive multi-document visual summarization system. In Machine Learning and Knowledge Discovery in Databases: European Conference, ECML PKDD 2011, Athens, Greece, September 5-9, 2011, Proceedings, Part III 22. Springer, 569–584.
- Zhang et al. (2023a) Zheng Zhang, Jie Gao, Ranjodh Singh Dhaliwal, and Toby Jia-Jun Li. 2023a. Visar: A human-ai argumentative writing assistant with visual programming and rapid draft prototyping. In Proceedings of the 36th Annual ACM Symposium on User Interface Software and Technology. 1–30.
- Zou and Hyland (2019) Hang Zou and Ken Hyland. 2019. Reworking research: Interactions in academic articles and blogs. Discourse studies 21, 6 (2019), 713–733.
Appendix A Prompts
A.1. Prompt for Initial LLM Draft for Baseline Tool
Figure 10 shows the prompt for the initial LLM draft for the baseline tool.
A.2. Prompts for Initial LLM Draft for Papers-to-Posts
A.3. Prompts for Working with Papers-to-Posts
Appendix B Additional Log Analysis Plots
Figures 25 and 26 show additional log analysis plots. Figure 27 shows the distribution of behavioral actions taken.
Appendix C Formative Study Technology Probe
Figure 28 shows the formative study technology probe.
Appendix D Updated Tool for Deployment Study
Figure 29 shows the updated tool for the deployment study.